OpenStudy (anonymous):

2x-5/x-3=4x+1/2x solve for x

OpenStudy (amistre64):

hard to make out the problem

OpenStudy (anonymous):

4x^2-10x=4x^2-12x+x-3 -10x=-11x-3 x=-3

OpenStudy (anonymous):

2x-5 ______ = 4x+1 x-3 ____ 2x

OpenStudy (amistre64):

that easier to read :)

OpenStudy (anonymous):

thanks roh...i got that too, guess i did it right lol

OpenStudy (amistre64):

its a techniqued refered to sometimes unmathematically as "cross multiply"

OpenStudy (amistre64):

since the end result looks like you simply multiplied the denominators across to the numerators

OpenStudy (anonymous):

lol yea....its just that i havent taken math in a year and now i have to take ap calc and im kinda rusty

OpenStudy (amistre64):

:) good luck with it

OpenStudy (anonymous):

thank you :)